# Deep Linear Hawkes Processes

### v1

@@ D.5 Model Calibration
Model Probabilistic Calibration Error (PCE) ($\downarrow$)
@@
RMTPP 13.67 7.71 3.51 9.97 1.78 11.54 4.23 13.07
NHP 8.16 0.14 1.27 6.66 1.42 4.60 5.87 7.42
SAHP 10.79 6.23 1.75 3.11 1.38 10.44 2.00 20.02
THP 12.38 5.60 3.33 16.32 1.92 10.97 1.05 13.73
AttNHP 6.50 1.08 0.53 2.85 1.46 1.36 4.84 OOM
IFTPP 1.54 23.35 0.47 0.61 0.68 0.30 1.89 15.14
DLHP (Ours) 5.81 0.91 0.15 1.91 0.82 0.73 10.35 12.16
@@
Model Expected Calibration Error (ECE) ($\downarrow$)
@@
RMTPP 6.42 2.06 2.24 1.52 2.09 2.49 2.87 8.22
NHP 8.02 0.41 0.77 5.28 1.15 2.73 1.92 4.21
SAHP 8.93 8.66 6.96 2.87 1.48 7.79 5.58 4.80
THP 2.14 1.27 1.65 6.51 0.94 4.24 2.42 9.95
AttNHP 2.35 0.33 0.57 2.56 1.31 0.51 1.97 OOM
IFTPP 0.60 2.93 0.87 1.48 1.41 0.59 1.66 1.92
DLHP (Ours) 1.30 0.45 0.52 1.48 1.74 0.31 1.31 2.47
@@
Finally, in Figs. 11 and 12 we plot the log-likelihood of time and mark respectively, versus their corresponding calibration results, to provide an overall view of the performances of different models. Our DLHP model consistently achieves higher log-likelihood while maintaining good calibration on both time and mark components on most datasets.

### v2

@@ D.7 Model Calibration
Model Probabilistic Calibration Error (PCE) ($\downarrow$) Avg. Ranking ($\downarrow$)
@@
RMTPP 13.67 (0.03) 7.93 (0.62) 3.50 (0.03) 0.22 (0.16) 1.94 (0.10) 1.56 (0.01) 3.63 (0.37) 12.60 (0.37) 6.1
SAHP 12.04 (1.02) 8.51 (1.86) 2.52 (0.99) 3.18 (0.21) 1.50 (0.57) 2.53 (1.86) 2.28 (0.44) 20.20 (1.09) 5.3
THP 12.38 (0.05) 5.68 (0.08) 3.34 (0.02) 6.36 (0.04) 2.06 (0.11) 1.02 (0.08) 1.10 (0.06) 13.46 (0.45) 5.4
IFTPP 1.59 (0.09) 23.85 (0.26) 0.40 (0.10) 1.61 (0.74) 0.84 (0.34) 0.46 (0.44) 1.75 (0.33) 16.58 (3.34) 2.6
MHP 12.22 (0.04) 4.89 (0.16) 3.43 (0.05) 8.77 (0.40) 1.58 (0.13) 1.25 (0.05) 6.21 (0.18) 15.24 (0.92) 6.0
NHP 8.45 (0.28) 0.20 (0.19) 0.87 (0.50) 7.40 (0.68) 1.51 (0.11) 4.70 (0.13) 5.92 (0.14) 7.70 (0.49) 3.6
AttNHP 6.36 (0.63) 2.09 (0.85) 0.84 (0.27) 3.08 (0.16) 1.65 (0.24) 1.43 (0.14) 4.70 (0.33) OOM 3.7
S2P2 (Ours) 5.88 (0.17) 0.44 (0.27) 0.55 (0.33) 2.07 (0.32) 1.03 (0.15) 1.38 (0.52) 11.70 (0.68) 12.06 (0.54) 2.8
@@
Model Expected Calibration Error (ECE) ($\downarrow$) Avg. Ranking ($\downarrow$)
@@
RMTPP 6.58 (0.15) 3.99 (4.28) 2.42 (0.16) 1.89 (0.24) 2.10 (0.27) 2.47 (0.45) 2.79 (0.43) 8.47 (0.31) 5.8
SAHP 8.17 (2.00) 6.27 (2.23) 6.77 (0.21) 2.68 (0.35) 1.71 (0.77) 6.26 (4.30) 5.41 (0.26) 5.85 (1.95) 6.8
THP 2.06 (0.17) 1.26 (0.11) 1.76 (0.07) 6.51 (0.03) 0.81 (0.14) 3.42 (0.70) 2.16 (0.39) 8.95 (0.91) 5.0
IFTPP 0.46 (0.10) 0.95 (1.12) 0.55 (0.19) 1.20 (0.20) 1.28 (0.54) 0.66 (0.05) 1.39 (0.23) 1.99 (0.61) 1.8
MHP 1.65 (0.16) 1.18 (0.12) 1.91 (0.11) 4.15 (0.36) 0.82 (0.18) 2.83 (0.50) 2.22 (0.24) 10.00 (1.71) 4.8
NHP 8.30 (0.21) 0.35 (0.06) 0.79 (0.10) 5.59 (0.69) 1.31 (0.16) 3.41 (0.41) 2.24 (0.32) 4.18 (0.69) 4.9
AttNHP 3.13 (0.61) 0.52 (0.16) 0.56 (0.10) 2.47 (0.12) 1.37 (0.42) 0.61 (0.16) 2.23 (0.50) OOM 3.4
S2P2 (Ours) 0.88 (0.34) 0.52 (0.13) 0.58 (0.12) 1.96 (0.67) 1.98 (0.19) 1.01 (0.63) 1.62 (0.24) 2.51 (0.44) 3.0
@@
Finally, in Figs.˜14 and 15 we plot the log-likelihood of time and mark respectively, versus their corresponding calibration results, to provide an overall view of the performances of different models. Our S2P2 model consistently achieves higher log-likelihood while maintaining good calibration on both time and mark components on most datasets.